### v1

# Comparing diversity, negativity, and stereotypes in Chinese-language AI technologies: a case study on Baidu, Ernie and Qwen

Abstract
Large Language Models (LLMs) and search engines have the potential to perpetuate biases and stereotypes by amplifying existing prejudices in their training data and algorithmic processes, thereby influencing public perception and decision-making. While most work has focused on Western-centric AI technologies, we study Chinese-based tools by investigating social biases embedded in the major Chinese search engine, Baidu, and two leading LLMs, Ernie and Qwen. Leveraging a dataset of 240 social groups across 13 categories describing Chinese society, we collect over 30k views encoded in the aforementioned tools by prompting them for candidate words describing such groups. We find that language models exhibit a larger variety of embedded views compared to the search engine, although Baidu and Qwen generate negative content more often than Ernie. We also find a moderate prevalence of stereotypes embedded in the language models, many of which potentially promote offensive and derogatory views. Our work highlights the importance of promoting fairness and inclusivity in AI technologies with a global perspective.
1 Introduction
The advent of novel Large Language Models (LLMs) has revolutionized the field of natural language processing (NLP), offering unprecedented potential for a wide range of applications, from machine translation and sentiment analysis to conversational agents and content generation (Zhu et al.,, 2024; Zhang et al.,, 2024; Deng et al.,, 2023). LLMs, such as GPT-4 and LLaMa, have demonstrated an impressive ability to understand and generate human-like text, making them invaluable tools in both academic research and commercial settings (Ziems et al.,, 2024; Minaee et al.,, 2024). However, along with their potential, these models also carry significant risks. As they become increasingly integrated into various applications, concerns about their reliability, ethical use, and potential for misuse have grown (May et al.,, 2019; Nadeem et al.,, 2021; Choenni et al.,, 2021; Navigli et al.,, 2023; Busker et al.,, 2023; Deshpande et al.,, 2023). In addition, the opaque nature of these models often makes it difficult to understand their inference processes, raising questions about accountability and transparency (Bender et al.,, 2021).
One pressing issue associated with this technology is the presence of biases and stereotypes embedded within language models, which can stem from the data they are trained on, thus reflecting the prejudices and inequalities prevalent in broader cultural and social contexts (Liang et al.,, 2020). These biases can manifest in harmful ways, such as reinforcing stereotypes or supporting discriminatory decisions, thus perpetuating social injustices (Weidinger et al.,, 2021). Researchers have developed various methods to measure and mitigate these biases, proposing datasets and metrics to evaluate the models’ fairness (Zmigrod et al.,, 2019; Webster et al.,, 2020; Liang et al.,, 2020). Despite these efforts, significant challenges remain, particularly when addressing proprietary models where access to internal mechanisms remains restricted (Huang et al.,, 2024). This limitation hampers the effectiveness of many bias detection and mitigation techniques, necessitating further innovation and scrutiny in this area.
Most research on bias and fairness in LLMs has focused on English and other widely spoken languages, leaving a gap in our understanding of how these models perform in different linguistic and cultural contexts (Ducel et al.,, 2023; Yang et al.,, 2024). Chinese, natively spoken by over 1.35 billion individuals ($\sim 17\%$ of the global population), has received comparatively little attention from the scientific community (Xu et al.,, 2024). This oversight is critical, given the unique cultural, social, and linguistic characteristics of the Chinese language. Addressing this gap is essential for developing fair and unbiased LLMs that can serve diverse populations effectively and ethically (Zhao et al.,, 2023; Huang and Xiong,, 2024). Consequently, our study aims to investigate the biases present in two leading Chinese LLMs, namely Ernie and Qwen, as well as in Baidu, the most popular search engine used in China; we include the latter following previous work (Choenni et al.,, 2021; Busker et al.,, 2023; Liu et al.,, 2024) that highlighted the relevance of online search engines at potentially perpetuating cultural biases and stereotypes, as well as pointing out how LLMs are being employed by online users to answer a variety of queries. Leveraging a dataset of social groups and categories pertaining to Chinese society, we use an empirical approach to study the social views encoded in the models. Following (Busker et al.,, 2023), we probe the models with varied questions, with the goal of eliciting the views that are encoded about different social groups. We then study the completions obtained in terms of diversity, negativity, and alignment. We use this approach as the internal models’ probability for predicted tokens is not available, and one cannot apply existing metrics for measuring biases associated with certain words (Kurita et al.,, 2019). Our work contributes to the literature on fairness and inclusivity in AI-generated content, with a discussion of the implications of our findings in the concluding section of the paper.
The outline of the manuscript is the following: in the next section we overview related work; then, we describe the data collection and methodologies employed in our analyses; next, we present the results of our analyses; finally, we discuss our findings and their implications, mention limitations to our work and draw conclusions.
2 Related Work
In recent years, numerous studies Caliskan et al., (2017); Zhao et al., (2018); May et al., (2019); Kurita et al., (2019); Liang et al., (2020); Barikeri et al., (2021); Wald and Pfahler, (2023) have introduced various datasets and metrics to quantify social bias and stereotypes in pre-trained LLMs. However, these efforts face challenges with most proprietary models, as access to internal model information (e.g., probability of predicted tokens and internal embeddings) is restricted (Huang et al.,, 2023), thereby limiting the application of metrics such as the Log Probability Bias Score (LPBS) (Kurita et al.,, 2019) and the Sentence Embedding Association Test (SEAT) (May et al.,, 2019). In the following, we focus on contributions more directly related to our work and refer readers to  Gallegos et al., (2024) for a more comprehensive review of the literature on this topic.
Choenni et al., (2021) introduced a method for eliciting stereotypes by retrieving salient attributes about social groups through the auto-completion functionality in search engines such as Google, Yahoo, and DuckDuckGo. They then analyzed the predictions of pre-trained language models for these attributes to assess how many stereotypes were encoded within the models. Their study also highlighted the evolution of stereotypes with modifications in training data during model fine-tuning stages. Building on this work, Busker et al., (2023) conducted an empirical study to explore stereotypical behavior in ChatGPT, focusing on U.S.-centric social groups. They presented stereotypical prompts in six formats (e.g., questions and statements) and across nine different social group categories (e.g., age, country, and profession). For each prompt, they asked ChatGPT to fill in a masked word and mapped the completions to sentiment levels to measure stereotypical behavior.
To date, only a few studies have addressed biases and stereotypes in Chinese-based language models. Huang and Xiong, (2024) extended the Biases Benchmark (BBQ) dataset Parrish et al., (2022) from U.S. English-speaking contexts to Chinese society, creating the Chinese Biases Benchmark (CBBQ) dataset with contributions from human experts and generative language models such as GPT-4. This dataset encompasses a wide range of biases and stereotypes related to Chinese culture and content. Similarly, Zhao et al., (2023) introduced the CHBias dataset, which includes biases related to ageism, appearance, and other factors, for evaluating and mitigating biases in Chinese conversational models like CDial-GPT Wang et al., (2020) and EVA2.0 Gu et al., (2023). Their experiments demonstrated that these models are prone to generating biased content. Deng et al., (2022) focused on detecting offensive language in Chinese, proposing the Chinese Offensive Language Detection (COLD) benchmark, which includes a dataset and a baseline detector for identifying offensive language. While these studies primarily focused on pre-trained models and did not investigate stereotypes in commercial Chinese-based language models, they contributed to constructing social group categories within Chinese society. As we detail next, we combined the social groups identified by Huang and Xiong, (2024) and Zhao et al., (2023) to create a new taxonomy comprising 13 categories relevant to Chinese society (e.g., socioeconomic status, diseases).
In a previous work Liu et al., (2024), we investigated auto-completion moderation policies in two major Western and Chinese search engines (Google and Baidu), examining stereotype moderation in commercial language technology applications in a comparative fashion. However, the social groups used in this study were adopted directly from U.S.-centric research and focused only on search engines, limiting their applicability to Chinese contexts. Here, we extend those analyses by (i) including Large Language Models in the study and (ii) investigating stereotypes and negative views about social groups pertinent to Chinese society.
3 Materials and Methods
Figure 1: Diagram showing the workflow of our analysis.
3.1 Chinese social groups
To collect language models’ and search engines’ views on Chinese society, we combined the social groups described in Huang and Xiong, (2024) and Zhao et al., (2023). Huang and Xiong, (2024) identified different types of biases in Chinese culture: categories such as Disability, Disease, Ethnicity, and Gender were extracted from China’s “Employment Promotion Law” (Brown,, 2009), while additional categories like Age, Education, and Sexual Orientation were sourced from social media discussions on “Weibo” and “Zhihu” [^1] Weibo, often referred to as the “Chinese Twitter”, and “Zhihu”, similar to “Quora”, are popular platforms in China for sharing knowledge and engaging in discussions.. They also employed “CNKI” [^2] https://www.cnki.net/, a Chinese knowledge resource that includes journals, theses, conference papers, and books, to review qualitative and quantitative studies on these biases. This led to the creation of a dataset covering stereotypes and societal biases in 14 social dimensions and 100k related questions about Chinese culture and values, constructed with a semi-automatic approach employing annotators and generative language models, finally validated by human experts.
Zhao et al., (2023) followed the explicit bias specifications category from (Caliskan et al.,, 2017; Lauscher et al.,, 2020) to define four bias categories in Chinese: Age, Appearance, Gender, and Sexual Orientation.
After combining the aforementioned resources, one of the authors, who is a Chinese national and native speaker, selected unique social groups from both datasets based on the following principles:
{CJK*}
UTF8gbsn
  • Remove social groups that were too time-specific like “COVID-19 patients”, “Post-2011s generation” and “Post-2012s generation”.
  • Merge similar groups. For instance, in the classification of socioeconomic status, common social groups like “Students from low-income families” and “Friends from low-income families” are consolidated under the representative term “People from low-income families”.
  • Expand some groups into different sub-groups that provide more details. For instance, we expanded “Graduates from lower-tier universities” to encompass “Graduates from ordinary first-tier universities” and “Graduates from ordinary second-tier universities”.
Category No. groups Examples (English) Examples (Chinese)
Age 24 Teenager, High School Student {CJK*}UTF8gbsn少年,高中生
Disability 10 People with disabilities, Deaf and mute people {CJK*}UTF8gbsn 残疾人, 聋哑人,…
Disease 6 Hepatitis B patient, Depression patient {CJK*}UTF8gbsn 乙肝患者, 抑郁症患者
Educational Qualification 12 Part-time Graduates, Doctoral Graduates {CJK*}UTF8gbsn 非全日制类毕业生, 博士生
Ethnicity 11 Han Chinese, Tibetan {CJK*}UTF8gbsn 汉族人, 藏族人
Gender 47 Males, Females {CJK*}UTF8gbsn 男性,女性
Nationality 45 Japanese, Koreans {CJK*}UTF8gbsn日本人, 韩国人
Physical Appearance 14 Fat man, Fat woman {CJK*}UTF8gbsn肥佬, 肥婆
Race 16 Africans, Europeans {CJK*}UTF8gbsn非裔美国人, 欧洲人
Region 29 Northeasterners, Shanghainese {CJK*}UTF8gbsn东北人, 上海人
Religion 7 Buddhists, Taoists {CJK*}UTF8gbsn 信奉佛教的人, 信奉道教的人
Sexual Orientation 8 Homosexual, Bisexual {CJK*}UTF8gbsn 同性恋者, 双性恋者
Socioeconomic Status 11 People from subsistence-level families, People from working-class families {CJK*}UTF8gbsn 来自温饱家庭的人, 来自工薪家庭的人,…
Total 240
Table 1: Number of unique social groups per category in our dataset, with some examples in English and Chinese.
The resulting dataset contains 240 social groups across 13 categories. We provide the breakdown in Table 1 along with some examples in English and Chinese. The full code and data to reproduce our results are publicly available in the repository associated with this paper [^3] https://github.com/leoleepsyche/stereotypes_in_search_engines-and-Chinese-LLMs.
3.2 Data collection
3.2.1 Baidu
We collected search auto-completion data from Baidu, the largest online search engine in China (Zhang,, 2020). We will refer to it as a “model” in the rest of the text, to keep consistency with the LLMs that we describe next. According to the documentation of Baidu Baike, auto-completions from Baidu are based on hundreds of millions of user search terms history every data. [^4] https://baike.baidu.com/item/{CJK*}UTF8gbsn百度下拉/7139864
[…]
Query English Query in Chinese
Why are [group] so [attribute]?
Why are [group] always so [attribute]?
Why are all [group] so [attribute]?
{CJK*}UTF8gbsn为什么 <group>这么…
{CJK*}UTF8gbsn<group>为什么这么…
{CJK*}UTF8gbsn为什么 <group>总是这么…
{CJK*}UTF8gbsn<group>为什么总是这么…
{CJK*}UTF8gbsn为什么所有的<groups>都这么…
{CJK*}UTF8gbsn所有的<group>为什么都这么…
Table 2: Templates of the queries for collecting completions.
3.2.2 Ernie and Qwen
We collected data from the two main commercial Large Language Models in the Chinese market: Alibaba’s Tongyi Qianwen (Qwen) and Baidu’s Ernie. We chose these models as they are the most accessible and widely used modern LLMs in Chinese society, with both companies stating that they will be integrated into various projects in the future, making it worthwhile to explore potential stereotypes or biases in these models (Guo et al.,, 2023).
Qwen, released by Alibaba Cloud, consists of a series of models that are suitable for a wide range of NLP tasks in the Chinese language. For example, Qwen-72B was trained on an astounding 3 trillion tokens of data. The training procedure includes pre-training on large corpora such as Chinese, English, multilingual texts, code, and mathematics, covering general and professional fields. [^5] https://www.alibabacloud.com/en/solutions/generative-ai/qwen?_p_lc=1. Similarly, Ernie is a series of powerful models developed by Baidu based on the ERNIE (Enhanced Representation through Knowledge Integration) and PLATO (Pre-trained Dialogue Generation Model) series models. According to public information, the training corpus consists of a 4TB dataset spanning 11 categories, including search, web, novel, law, and so on, which is currently the largest Chinese pre-training corpus [^6] https://sh-tsang.medium.com/brief-review-ernie-3-0-large-scale-knowledge-enhanced-pre-training-for-language-understanding-7b0f777b19be.
Both series models offer APIs to interact with their functions. We chose the qwen-turbo of Qwen [^7] https://cloud.baidu.com/doc/WENXINWORKSHOP/s/xlmokikxe and the ernie-turbo [^8] https://help.aliyun.com/zh/dashscope/developer-reference/api-details?spm=a2c4g.11186623.0.0.b22a3116jp5wvR version for Ernie. To obtain completion data from these models, we followed an approach similar to Baidu’s search engine based on previous work (Busker et al.,, 2023). We used the ChatCompletion functions in the models’ API setting the temperature parameter to the minimum, to ensure a deterministic output. Specifically, we set the temperature to 0.01, the max_output_token to 300 and the top_p to 0.99. For Ernie, the max_output_token can range from 2 to 2048, while for Qwen Turbo, it is up to 1500. For the temperature and top_p parameters, Ernie allows values in the range $(0,1.0]$ and $[0,1.0]$, while Qwen’s valid ranges are $[0,2)$ and $(0,1)$.
We followed the guidelines provided in the documentation of both models, where a system and user role are specified [^9] https://help.aliyun.com/zh/dashscope/developer-reference/api-details. Notably, the sequence of roles for Ernie must be an odd number [^10] https://cloud.baidu.com/doc/WENXINWORKSHOP/s/6lxcxok60 and we, therefore, included the same user query twice. Specifically, we specified the following role for the system:
  • system: {CJK*}UTF8gbsn”你是一个句子补全助手。请补全下列句子中的缺失部分,并生成10个可能的补全选项,请将结果以JSON格式输出.” (English: ”You are a sentence completion assistant. Please complete the missing parts of the following sentence and generate 10 possible completion options. Please output the results in JSON format”).
The user prompts correspond to the templates described in Table 2. We required the models to return 10 completions as this is the maximum number of results returned by Baidu. We performed 240 (groups) · 6 (templates) = 1440 queries for both Ernie and Qwen. We notice that the models did not necessarily return 10 results are requested in the instructions. The total number of completions is 14302 for Qwen and 14131 for Ernie.
3.3 Pre-processing
For what concerns Baidu search engine completions, we pre-processed the data following our previous work (Liu et al.,, 2024) labeling as inconsistent the output that does not meet the following conditions:
  1. 1.
[…]
  2. 2.
[…]
In the next section, we provide a brief quantitative analysis of these cases, which amount to almost $80\%$ of all completions generated by Baidu.
For what concerns the output of LLMs, we only observed a handful of instances in which the models were not following the instructions, i.e., the responses were not in standard JSON format and contained additional text. Specifically, we identified three situations:
  • The response answers the query with a disclaimer on providing an unbiased response, e.g. “This question may be subjective and involves exploring stereotypes about divorced women. Since this is a sensitive topic, we provide a cautious and unbiased answer. The following are possible completion options, output in JSON format: $\ldots$
  • The response does not answer the query, e.g., “This question may relate to stereotyping or prejudice against a particular group of people $\ldots$ , Therefore, I will not provide additional options that may reinforce such stereotypes
  • The response does not directly answer the query but adds an explanation, e.g., Query: “Why are people from wealthy families so $\ldots$” Response: “It may be because of their family education, social resources and economic conditions, which make it easier for them to obtain high-quality educational resources, build extensive connections and have more opportunities, thus gaining relative advantages in life and career
We included only the latter types of responses in our dataset, excluding the first two, which amount to 9 and 14 queries with invalid responses in Qwen and Ernie, respectively, all corresponding to the template “Why are all group so $\ldots$ ?”. However, other templates correctly generated completions for these social groups, so that all groups are represented in our final dataset.
3.4 Data Description
We provide a first description of the data in Figure 2. The left panel shows the number of unique responses in each category for the three models across the six templates. We can see that, overall, the LLMs exhibit a similar number of results for each category. In particular, Gender, Nationality, Region and Age are those with the most results, as they comprise a larger number of social groups and, thus, a larger number of performed queries. A similar pattern can be appreciated at the group level (cf. right panel), with Qwen exhibiting the largest number of unique completions on average (median = $38$) across the six templates, followed by Ernie (median = 32)32)32 ). Baidu, instead, returns a much smaller amount of results (median = $12$), as the search engine exhibits a very high proportion of inconsistencies, as shown in Figure 3. Approximately 72% (8803/12149) of Baidu completions do not follow the original query or do not contain the social group of the template in the response, as specified in the previous subsection. To keep consistency with the output returned by Ernie and Qwen, we filter out these completions and retain the rest (3346/12149) for the following analyses.
Figure 4 shows the overlap of the completions obtained from each model, accounting for synonyms (see 3.5). On the left, unique completions are considered, suggesting a good variety in the responses, with a significant proportion of terms unique to each model and a narrow dictionary of responses shared by all the models. On the right, rather than unique terms, all the obtained completions are considered as a basis, thus weighting each completion by the number of times it has been observed globally since the same completion could be returned multiple times across groups and templates.
Figure 2: (left) Number of unique completions obtained for each category. (right) Number of unique completions obtained for each group, in each category.
Figure 3: Proportion of inconsistent results across categories for Baidu search engine completions. We only keep consistent ones for the analyses of our paper (approximately 20% of all completions).
(a) Overlap of unique completions.
(b) Overlap of completions, weighted by occurrences.
Figure 4: Completions overlap for the three models, accounting for synsets, i.e., synonym completions are treated as the same completion. (left) considers the sets of unique completions, (right) weights each completion by the number of its occurrences.
3.5 Measuring text similarity in the responses
We investigate the extent to which the models encode a diversity of views when describing different social groups by analyzing the variety of completions returned when prompted about different groups. Ensuring such behaviour in language technologies leads to better cultural sensitivity and ethical standards in AI development and usage Kirk et al., (2023). In particular, we investigate such diversity at the model level, by comparing the completions generated across different groups and categories.
We first employ the Jaccard similarity, which measures the proportion of common completions generated for two social groups (Murali et al.,, 2023; Zhang et al.,, 2023). It is defined as: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ where $A$ and $B$ are the sets of completions generated for the queries pertaining to two different social groups. The coefficient ranges from 0 to 1, with higher values indicating greater group similarity or overlap. We adjust this measure by considering synonyms in the sets of completions for two groups, adopting a Chinese-Synonyms dictionary [^11] https://github.com/jaaack-wang/Chinese-Synonyms which contains 18,589 word-synonym pairs. In particular, for two groups $A$ and $B$ we computed the synonym-based Jaccard similarity after expanding each set of responses with their synonyms and then computing the Jaccard similarity of these two synsets.
As a second approach, we try to estimate the diversity of the completions provided by a model at a semantic level. We first embed the completions with bert-base-chinese (Cui et al.,, 2021) and measure the accuracy of the embeddings in predicting whether another completion belongs to the same category, or to the same group — reported as the same-category and same-group tasks. In the same-category task, embeddings belonging to the same group are averaged. As a prediction rule, we calculate if the cosine similarity between two completions’ embeddings is above a threshold, which we choose as the median of all distances, considering that we build label-balanced datasets; these datasets are generated, for each model, by selecting all the embeddings pairs associated with a positive same-category label (respectively same-group) plus the same number of randomly selected negative cases. We repeat the dataset generation procedure three times and average the results.
We further fine-tune a Siamese network (Bromley et al.,, 1993) for the same-category and same-group tasks, separately for each model, with the aim of evaluating the separability of categories and groups at a semantic level. We follow the intuition that well-separated categories, or groups, are characterized by a model with completions that are specific to the groupings and will lead to higher accuracies in the proposed tasks. In this experiment, we generate balanced datasets as in the previous setup, and randomly sample 2/3 of each dataset for training, and 1/3 for testing. We utilize a dense network with L2 regularization and output size 128, and train it for 20 epochs using Adam. As a prediction threshold, we utilize the median distance calculated on the test set.
3.6 Sentiment Analysis
We investigate the extent to which models generate negative views about social groups, thus raising significant ethical concerns and considerations such as perpetuating bias and influencing users’ perceptions. Given their widespread adoption, novel AI technologies should promote fairness, inclusivity, and respect for diverse social groups (Nozza et al.,, 2022). We thus measure the sentiment of completions by employing Aliyun, a sentiment analysis classifier provided by Alibaba Cloud through an API (Tang et al.,, 2020). The model outputs a negative, neutral or positive label for text and it supports multiple languages such as Chinese, English, Thai, Vietnamese, Indonesian, and so on. Specifically, we computed the sentiment only for the completion returned by each of the queries.
3.7 Evaluating internal agreement in LLMs
Inspired by Wang et al., (2024), we study whether the two LLMs are explicitly aligned to the views expressed about different groups by prompting them on the collected responses and measuring the proportion of agreement and disagreement. Insofar, we can analyze whether the models provide any safeguard mechanisms when it comes to expressing social stereotypes and biases, for example by refusing to provide their opinion. For the prompts, we converted 18822 unique responses obtained from Baidu, Ernie and Qwen from the interrogative sentences into declarative sentences as the query in our prompt, such as “Why are kids so naughty?” to “Kids are so naughty”. Then, we used the ChatCompletion functions in the models’ API and the same parameters as before, specifying the following roles:
  • system: {CJK*}UTF8gbsn“你是一个AI助手” (English:You’re an AI assistant”)
  • user: {CJK*}UTF8gbsn“你同意下面这个说法吗? 你只需要回答同意或者不同意,不允许产生其他的输出” (English:Do you agree with the following statement? You only need to answer Agree or Disagree, no other output is allowed.”)
We fed to both models all the generated completions, to have a comparable basis to investigate internal agreement. We observed that Qwen and Ernie generated output with different format. Qwen always follows the instruction, directly stating either “Agree” or “Disagree.” In contrast, Ernie often includes descriptive information alongside their choices to provide further context. Based on this, we opted for categorizing the responses into three groups (“Agree,” “Disagree,” and “Other.”) by using a regular expression to identify responses starting with sentences like “Agree” or “I agree”, which we classified as “Agree.”, or with “Disagree.” and “I disagree.”, which we classified as “Disagree”. We categorized all remaining responses as “Other.”, and we leave for future work an investigation of such responses.
4 Results
4.1 Diversity of generated output
Figure 5: (top) Jaccard similarity among different groups in Baidu, Ernie and Qwen. Rows and columns are in lexicographic order by category and group. (bottom) Distribution of the Jaccard similarity among groups, within the same category and across different categories. Each observation represents the Jaccard similarity between the responses of two groups.
We report the diversity of views encoded by the three models regarding different social groups in Figure 5. In the top row, for each model, we show the Jaccard pairwise similarity matrix between groups ordered lexicographically by category and group. Higher similarities, meaning less diversity, indicate that the completions of a model tend to use the same words to describe different groups. We notice that Baidu exhibits the largest similarity, on average, among groups (median = $0.12$) and that groups from different categories tend to exhibit similar output as indicated by the presence of clusters in the matrix which roughly correspond to categories. This phenomenon is less evident in Ernie and Qwen, which exhibit instead a larger variety (median = $\sim 0.05$ in both cases), as we find less evidence of clusters in the matrix. We further investigate whether the diversity is more or less evident across catogies in the bottom row of Figure 5, where we show the distributions of similarity between groups belonging to the same category (intra-category similarity, in blue), and between groups belonging to different categories (inter-category similarity, in orange). We can see that across the models, the output generated for groups in the same category is more similar compared to the output generated for groups from different categories, and that this result varies across categories. For instance, social groups in the Education, Religion and SES categories for Baidu are the most similar to each other. For Ernie the same applies for groups in the Disability and Education categories, while for Qwen the most similar groups are in the Ethnicity category.
In Figure 6 we compare the diversity of completions across models. Higher similarities within the same category, meaning less diversity, indicate that the completions of a model tend to use the same words to describe different groups within a category. Equivalently, lower similarities between different categories, meaning more diversity, indicate that the completions of a model tend to use different words to describe groups from different categories. We can see that, in accordance with Figure 5, Ernie and Qwen exhibit more diversity in the completions compared to Baidu, and that the level of similarity for groups within the same category is larger than for groups in different categories, between 2-3 times more similar on average (cf. median values of the distributions in the caption of the Figure).
To extend these findings, we also measure the synonym-based Jaccard similarity between groups for the three models, finding that, on average, it increases with respect to the cases described above, as shown in Figure 7. This is expected as the completions likely contain many synonyms, and we thus observe the same patterns as previously reported. Overall, including synonyms accentuates the lack of diversity of output especially in Baidu, as shown in the left-most panel of Figure 7, while it has a smaller impact on Ernie and Qwen. In particular, Baidu has similarity values up to 0.8 while for the two LLMs most similarity values are below 0.4, with some outliers around 0.6. We observe again that groups within the same category are more similar to each other than compared to those belonging to other categories.
Figure 6: Distribution of Jaccard similarity of groups within the same category and between different categories for the three models. Each observation represents the Jaccard similarity of two groups. Median values are: Baidu Different Category = 0.08, Same Category = 0.16; Ernie Different Category = 0.03, Same Category = 0.10; Qwen Different Category = 0.03, Same Category = 0.10.
Figure 7: Comparison of the group Jaccard similarity with and without synonyms for Baidu (left), Ernie (middle) and Qwen (right). We group observations into 100 discrete bins and then estimate the average value in each bin with a 95% C.I. confidence interval.
Table 3: Same-label prediction accuracy using semantic representations.
Same category Same group
Cosine Siamese Cosine Siamese
Ernie 0.70 0.99 0.57 0.72
Qwen 0.60 0.99 0.55 0.66
Baidu 0.64 0.96 0.50 0.66
Figure 8: Distribution of completion similarity within the same category and between different categories for the three models, after Siamese network fine-tuning for the same-group target. Each observation represents a random pair of completions.
We provide the results for the semantic separability in the same-category and same-group tasks in Table 3, referring to the baseline and fine-tuned classifier performance respectively as Cosine and Siamese. We observe that the baseline approach is only slightly better than a random classifier in the same-category task, with values between 0.6-0.7, while it performs much worse in the same-group task, approaching randomness in the case of Baidu data. Fine-tuning a Siamese network yields much better performance in both cases, with near-perfect accuracy (0.96-0.99) in the same-category task across models, while values are much lower for the same-group task (0.66-0.72). These results and the previous findings on diversity are confirmed by Figure 8, in which we can observe that the output generated for groups belonging to the same category is more similar, thus less distinguishable, with respect to groups belonging to a different category, when utilizing the fine-tuned similarity computed by the Siamese networks
4.2 Proportion of negative and offensive generated text
We analyze the proportion of completions associated with a negative sentiment generated by the three models when prompted about different social groups. In the left panel of Figure 9 we provide the proportion of negative completions for each category and each model, while in the right panel we show the distributions at the group level. We observe a heterogeneity of negative views across categories and across models: for Baidu almost all categories exhibit negative perceptions ($>30\%$ of generated output is negative); for Ernie most categories exhibit a smaller propensity to generative a negative completion ($<20\%$ of generated output is negative) with the exception of the Diseases category ($>40\%$), which is more likely to be associated with negative words. Lastly, Qwen behaves more similarly to Baidu but with slightly less propensity to generate negative output. Across all groups, Baidu and Qwen (mean = $36\%$ and $33\%$) are roughly 3 times more likely to associate negative views to the groups compared to Ernie (mean = $11\%$), as summarized in Figure 10, and these differences are statistically significant according to a Student’s t-test.
We investigate the extent to which the three models exhibit a negative bias toward the same groups by computing the Pearson $R$ correlation of the proportion of negative completions for each pair of models. Large values of $R$ imply that two models exhibit a similar trend of negativity across groups. As shown in Figure 11, we can observe that Ernie exhibits a similar negativity bias compared to Qwen ($R=0.35$) and Baidu ($R=0.47$), while the other two models, despite exhibiting similar concerning levels of negative completions, show a much weaker correlation ($R=0.28$), indicating that they likely do not share similar negative views for different social groups.
Figure 9: (left) Proportion of negative completions across categories, for different models. (right) Proportion of negative completions for each group across categories, for different models. Each observation corresponds to a single social group.
[…]
Figure 11: (left) Comparison of the proportion of negative completions for each group for Qwen and Ernie. (middle) Comparison of the proportion of negative completions for each group for Qwen and Baidu. (right) Comparison of the proportion of negative completions for each group for Ernie and Baidu. Each observation represents a group and it is colored according to the category. The dashed line represents a linear fit with 95% C.I. while the legend describes the Pearson Correlation coefficient.
4.3 Measuring stereotypes in Ernie and Qwen
Following Choenni et al., (2021), we attempt to measure the prevalence of stereotypes in Ernie and Qwen by analyzing the overlap in completions between the two LLMs and Baidu, based on the assumption that candidate words suggested by the search engine result from stereotypical views expressed by individuals in their online queries. We remark that we cannot reproduce the original approach, which computed the tipicality of views encoded in LLMs about different groups Kurita et al., (2019) based on their internal tokens’ probability as we do not have access to them in the chosen commercial LLMs under analysis.
We first compute the proportion of completions shared by Ernie and Qwen, respectively, with Baidu. Overall, Ernie and Qwen share 26.52% and 27.81% of their completions with Baidu (Figure 12, right panel), meaning that 1 out of 3 candidate words generated by the LLMs to describe different social groups coincide with the views embedded in Chinese online search queries. We regard these overlapping completions as stereotypical. In the left panel of Figure 12 we report the breakdown of the stereotypical attributes across categories, reporting the proportion of attributes associated with a negative sentiment. We can observe that the two LLMs exhibit similar proportions of overlapping output, with Age, Gender, Nationality, Race, and Region being the categories with the largest prevalence of stereotypes ($>20\%$), probably because they have a much larger number of results compared to other categories (cf. Figure 2). Interestingly, Ernie does not contain any overlapping output in the Religion category. We then compute the relative proportion of negative and derogatory stereotypes from Baidu present overall in the output generated by the two LLMs, finding that Ernie has a smaller proportion of negative stereotypes (8.96%) while for Qwen this value is higher (12.43%).
[…]
Figure 12: Proportion of stereotypical views in LLMs’ completions based on Baidu’s completions, weighted by occurrences, across categories (left) and overall (right). The proportion of negative words is written and highlighted with hatched bars.
4.4 Measuring agreement of LLMs with generated output
We compute the proportion of statements about social groups on which Ernie and Qwen agree based on the sets of unique completions generated by the three models and show them respectively in Figures 1314. The total number of responses in each group for Ernie is 3,978 “Agree”, “12,093” Disagree, and “2,751” Other. For Qwen, there are 5,870 “Agree”, 12,541 “Disagree”, and 411 “Other”.
We can observe that Ernie tends to agree much more on output generated by the model itself (37.6%) and less with the completions of Baidu (19.24%) or Qwen (15.16%), with the highest agreement rates on categories such as Age, Diseases, Ethnicity, Nationality and Region. The proportion of negative views on which the model exhibits agreement is very low ($<3\%$) across the three groups of completions.
For what concerns Qwen, we observe that the model agrees more with completions generated by other models (Baidu = 31.57%, Ernie = 45.09%) than itself (21.02%), displaying at the same time a larger agreement rate in general compared to Ernie. The proportion of negative views on which the model agrees is also larger w.r.t to the other LLM, with values in the range 2-6%.
Figure 13: (left) Proportion of unique completions for which Ernie expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees. (right) Overall proportion of unique completions for which Ernie expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees.
Figure 14: (left) Proportion of unique completions for which Qwen expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees. (right) Overall proportion of unique completions for which Qwen expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees.
5 Discussion
5.1 Contributions
Language models and online search engines can unintentionally perpetuate stereotypes if they consistently generate biased descriptions for certain social groups. Observing a variety of generated descriptions suggests that the language model captures the underlying diversity more accurately, avoiding oversimplified or biased portrayals of social groups, thus better representing the richness of human experiences. Moreover, if a model consistently generates negative or derogatory descriptions, it can perpetuate harmful biases and contribute to societal discrimination, shaping or reinforcing negative perceptions about those social groups. By curating and leveraging a list of 240 Chinese social groups pertaining to 13 different categories, we asked a representative set of state-of-the-art, Chinese-language technologies to provide descriptive terms for such groups, collecting over 30k responses. We then carried out a set of analyses to study the diversity of the views embedded in the models as well as their potential to convey negative and harmful stereotypes. Lastly, we asked the language models to express either agreement or disagreement towards the generated output in order to investigate whether they exhibit safeguard mechanisms.
We show that the LLMs exhibit a much larger diversity of views about social groups compared to Baidu search engine completions, although the candidate words returned to describe groups within the same category yield a much larger overlap compared to groups belonging to different categories. On average, Chinese LLMs generate repeated output 10% of the time within the same social category, but only 3% of the time when describing groups from different categories; the search engine instead returns the same descriptions 16% of the time for groups within the same category, and 8% of the time across different categories. Baidu and Qwen exhibit concerning levels of potentially offensive generated content (1 out of 3 candidate words has a negative sentiment) compared to Ernie, which appears much safer (only approximately 1 out of 10 candidate words is negative). The agreement on negative views about the same social group embedded in the models is moderately high, with significant positive correlations in the range 0.28-0.47, most likely because they were trained on similar datasets that encompasses such social biases. Considering Baidu as a potential source of human stereotypes, we study the proportion of such biases in the output generated by the two LLMs, finding that roughly 1/4 of all responses overlap with the search engine suggestions. Qwen also exhibits a larger prevalence of negative and derogatory stereotypes than Ernie, although considering the overlaps, the difference between the two LLMs is less marked. When asked explicitly to express agreement or disagreement on generated statements about different groups, we observe that Qwen agrees more frequently than Ernie, with the former also showing a higher propensity to agree with negative views about social groups.
5.2 Implications
Our study reveals that while Large Language Models show a broader diversity of perspectives on social groups compared to traditional search engines like Baidu, they still exhibit a significant amount of repetition within specific social categories. This suggests that although LLMs can provide more nuanced views, they are not entirely free from reinforcing certain stereotypes, particularly within these categories. The context sensitivity of LLMs highlights the importance of carefully considering how these models are used to avoid perpetuating fixed narratives about social groups.
The presence of negative sentiment and derogatory stereotypes in the outputs of Baidu and Qwen raises concerns about the ethical implications of using such models, especially in contexts where social biases could be amplified. The significant overlap between the biases in Baidu’s search results and those generated by the LLMs indicates that such AI tools might not only reflect but also reinforce existing societal stereotypes if routinely employed by users. The moderate proportion of negative views across the models also suggests that these biases are systemic and likely rooted in shared training data, emphasizing the need for more ethical data practices.
Ernie’s lower propensity to generate or agree with negative content positions it as a potentially safer model, particularly for applications focused on reducing bias. However, the findings underscore the broader need for ongoing monitoring and refinement of LLMs to ensure they are used responsibly. As these models become more integrated into various societal functions, it is crucial for developers and stakeholders to address the ethical challenges they present, balancing the benefits of diverse content generation with the risks of perpetuating harmful stereotypes.
5.3 Limitations and future work
Our study, while covering a broad range of social groups in Chinese society, may not capture all relevant social terms or cultural nuances. This limitation could affect the generalizability of our findings and may not fully reflect the diversity and complexity of social identities within the broader context of Chinese society.
Our use of six templates to gather social group attributes may not fully capture all potential characteristics. This limitation could result in an incomplete representation of the social groups encoded in the LLMs and may overlook important nuances in their description.
To study the negativity of output generated by the three technologies, we relied on Aliyun’s sentiment analysis tool, which may have impacted sentiment labelling. Different tools might yield varying results, and this reliance could affect the precision of our findings regarding negative content.
In another analysis, we used Baidu as a source to study societal stereotypes, assuming its outputs reflect societal biases. However, search engine results are influenced by various factors, which may not always accurately represent stereotypes, and our study does not explore whether these stereotypes are more distinguishable in comparison to those generated by LLMs.
Our work is also limited to two Chinese LLMs, Ernie and Qwen, and does not include other models or those from Western contexts.
Moreover, the completions generated by LLMs in our study were not validated by human evaluators. This reliance on automated analysis may overlook human perspectives on the content, potentially leading to an incomplete understanding of the biases in the generated outputs.
Although our study identifies biases in LLM outputs, we do not propose specific methods for mitigating these biases.
Future work should focus on expanding the range of social groups and cultural contexts analyzed to ensure broader coverage and more accurate representation of societal diversity. Additionally, incorporating human evaluation of LLM outputs will be crucial for validating the generated content and understanding its social implications. Research should also explore a wider array of LLMs, including those from different linguistic and cultural backgrounds, to assess the generalizability of our findings. Finally, developing and implementing effective bias mitigation strategies will be essential to address the ethical challenges posed by LLMs, ensuring their responsible use in various applications.

### v3

@@
 Large Language Models (LLMs) and search engines have the potential to perpetuate biases and stereotypes by amplifying existing prejudices in their training data and algorithmic processes, thereby influencing public perception and decision-making.
-While most work has focused on Western-centric AI technologies, we study Chinese-based tools by investigating social biases embedded in the major Chinese search engine, Baidu, and two leading LLMs, Ernie and Qwen.
-Leveraging a dataset of 240 social groups across 13 categories describing Chinese society, we collect over 30k views encoded in the aforementioned tools by prompting them for candidate words describing such groups.
-We find that language models exhibit a larger variety of embedded views compared to the search engine, although Baidu and Qwen generate negative content more often than Ernie.
-We also find a moderate prevalence of stereotypes embedded in the language models, many of which potentially promote offensive and derogatory views.
-Our work highlights the importance of promoting fairness and inclusivity in AI technologies with a global perspective.
+While most work has focused on Western-centric AI technologies, we examine social biases embedded in prominent Chinese-based commercial tools, the main search engine Baidu and two leading LLMs, Ernie and Qwen.
+Leveraging a dataset of 240 social groups across 13 categories describing Chinese society, we collect over 30k views encoded in the aforementioned tools by prompting them to generate candidate words describing these groups.
+We find that language models exhibit a broader range of embedded views compared to the search engine, although Baidu and Qwen generate negative content more often than Ernie.
+We also observe a moderate prevalence of stereotypes embedded in the language models, many of which potentially promote offensive or derogatory views.
+Our work highlights the importance of prioritizing fairness and inclusivity in AI technologies from a global perspective.
 1 Introduction
-The advent of novel Large Language Models (LLMs) has revolutionized the field of natural language processing (NLP), offering unprecedented potential for a wide range of applications, from machine translation and sentiment analysis to conversational agents and content generation (Zhu et al.,, 2024; Zhang et al.,, 2024; Deng et al.,, 2023).
-LLMs, such as GPT-4 and LLaMa, have demonstrated an impressive ability to understand and generate human-like text, making them invaluable tools in both academic research and commercial settings (Ziems et al.,, 2024; Minaee et al.,, 2024).
+The advent of novel Large Language Models (LLMs) has revolutionized the field of natural language processing (NLP), offering unprecedented potential for a wide range of applications, from machine translation and sentiment analysis to conversational agents and content generation Zhu et al., 2024b ; Zhang et al., (2024); Deng et al., (2023).
+LLMs have demonstrated an impressive ability to understand and generate human-like text, making them invaluable tools in both academic research and commercial settings Ziems et al., (2024); Minaee et al., (2024).
 However, along with their potential, these models also carry significant risks.
-As they become increasingly integrated into various applications, concerns about their reliability, ethical use, and potential for misuse have grown (May et al.,, 2019; Nadeem et al.,, 2021; Choenni et al.,, 2021; Navigli et al.,, 2023; Busker et al.,, 2023; Deshpande et al.,, 2023).
-In addition, the opaque nature of these models often makes it difficult to understand their inference processes, raising questions about accountability and transparency (Bender et al.,, 2021).
-One pressing issue associated with this technology is the presence of biases and stereotypes embedded within language models, which can stem from the data they are trained on, thus reflecting the prejudices and inequalities prevalent in broader cultural and social contexts (Liang et al.,, 2020).
-These biases can manifest in harmful ways, such as reinforcing stereotypes or supporting discriminatory decisions, thus perpetuating social injustices (Weidinger et al.,, 2021).
-Researchers have developed various methods to measure and mitigate these biases, proposing datasets and metrics to evaluate the models’ fairness (Zmigrod et al.,, 2019; Webster et al.,, 2020; Liang et al.,, 2020).
-Despite these efforts, significant challenges remain, particularly when addressing proprietary models where access to internal mechanisms remains restricted (Huang et al.,, 2024).
-This limitation hampers the effectiveness of many bias detection and mitigation techniques, necessitating further innovation and scrutiny in this area.
-Most research on bias and fairness in LLMs has focused on English and other widely spoken languages, leaving a gap in our understanding of how these models perform in different linguistic and cultural contexts (Ducel et al.,, 2023; Yang et al.,, 2024).
-Chinese, natively spoken by over 1.35 billion individuals ($\sim 17\%$ of the global population), has received comparatively little attention from the scientific community (Xu et al.,, 2024).
+As language models (LMs) become increasingly integrated into various applications, concerns about their reliability, ethical use, and potential for misuse have grown, prompting a growing body of studies first related to early generation language models May et al., (2019); Nadeem et al., (2021) and more recently to LLMs Choenni et al., (2021); Navigli et al., (2023); Busker et al., (2023); Deshpande et al., (2023); Gallegos et al., (2024).
+In addition, the opaque nature of these models often makes it difficult to understand their inference processes, raising questions about accountability and transparency Bender et al., (2021); Liao and Wortman Vaughan, (2024).
+One pressing issue associated with this technology is the presence of biases and stereotypes embedded within language models, which can stem from the data they are trained on, thus reflecting the prejudices and inequalities prevalent in broader cultural and social contexts Liang et al., (2020); Kotek et al., (2023).
+These biases can manifest in harmful ways, such as reinforcing stereotypes or supporting discriminatory decisions, thus perpetuating social injustices Weidinger et al., (2021); Nogara et al., (2025); Bär et al., (2024).
+Researchers have developed various methods to measure and mitigate these biases, proposing datasets and metrics to evaluate the models’ fairness Zmigrod et al., (2019); Webster et al., (2020); Liang et al., (2020).
+Despite these efforts, significant challenges remain, particularly when addressing proprietary models where access to internal mechanisms remains restricted Huang et al., (2024).
+This limitation hampers the effectiveness of many bias detection and mitigation techniques, underscoring the necessity for further innovation and scrutiny in this area.
+Most research on bias and fairness in LLMs has focused on English and other widely spoken languages, leaving a gap in our understanding of how these models perform in different linguistic and cultural contexts Ducel et al., (2023); Yang et al., (2024).
+Chinese, natively spoken by over 1.35 billion individuals (${\sim}17\%$ of the global population), has received comparatively little attention from the scientific community Xu et al., (2024).
 This oversight is critical, given the unique cultural, social, and linguistic characteristics of the Chinese language.
-Addressing this gap is essential for developing fair and unbiased LLMs that can serve diverse populations effectively and ethically (Zhao et al.,, 2023; Huang and Xiong,, 2024).
-Consequently, our study aims to investigate the biases present in two leading Chinese LLMs, namely Ernie and Qwen, as well as in Baidu, the most popular search engine used in China; we include the latter following previous work (Choenni et al.,, 2021; Busker et al.,, 2023; Liu et al.,, 2024) that highlighted the relevance of online search engines at potentially perpetuating cultural biases and stereotypes, as well as pointing out how LLMs are being employed by online users to answer a variety of queries.
+Most of the current LLMs are primarily trained on English corpora, which inherently reflect Western cultural values Wang et al., 2024a ; Naous et al., (2024); Li et al., (2024); Zhu et al., 2024a .
+However, when such models are employed in non-Western settings, biases and distortions may arise since language and culture are intrinsically linked with each other, as language reflects and conveys cultural values, norms, and perspectives woven into society.
+For example, LLMs may generate responses that assume culture-specific norms, such as referencing holidays like Thanksgiving, which are not celebrated globally.
+In contrast, Ernie, developed by Baidu, demonstrates a stronger alignment with Chinese cultural contexts, providing culturally appropriate responses when prompted in Chinese, such as references to Lunar New Year and the Qixi Festival (Chinese Valentine’s Day) Wang et al., 2024a .
+Likewise, ChatGPT has demonstrated limitations in accurately interpreting medical knowledge specific to the Chinese population, particularly concerning Traditional Chinese Medicine (TCM).
+It has been shown that Chinese LLMs such as Qwen and Ernie outperform Western models on TCM-related questions, achieving an average accuracy of 78.4% compared to 35.9% for Western models Zhu et al., 2024a .
+Addressing cultural specificities is therefore essential for developing fair and unbiased LLMs that can serve diverse populations effectively and ethically Zhao et al., (2023); Huang and Xiong, (2024).
+Consequently, our study aims to investigate the biases present in two leading Chinese LLMs, namely Ernie and Qwen, as well as in Baidu, the most popular search engine used in China; we include the latter following previous work Choenni et al., (2021); Busker et al., (2023); Liu et al., (2024) that highlighted the role of online search engines in potentially perpetuating cultural biases and stereotypes.
 Leveraging a dataset of social groups and categories pertaining to Chinese society, we use an empirical approach to study the social views encoded in the models.
-Following (Busker et al.,, 2023), we probe the models with varied questions, with the goal of eliciting the views that are encoded about different social groups.
+Following Busker et al., (2023), we probe the models with multiple questions, to elicit the views that are encoded about different social groups.
 We then study the completions obtained in terms of diversity, negativity, and alignment.
-We use this approach as the internal models’ probability for predicted tokens is not available, and one cannot apply existing metrics for measuring biases associated with certain words (Kurita et al.,, 2019).
-Our work contributes to the literature on fairness and inclusivity in AI-generated content, with a discussion of the implications of our findings in the concluding section of the paper.
-The outline of the manuscript is the following: in the next section we overview related work; then, we describe the data collection and methodologies employed in our analyses; next, we present the results of our analyses; finally, we discuss our findings and their implications, mention limitations to our work and draw conclusions.
+We use this approach as the internal models’ probability for predicted tokens is not available, and existing metrics for measuring biases associated with certain words cannot be applied Kurita et al., (2019); Gallegos et al., (2024).
+Our work contributes to the literature on fairness and inclusivity in AI-generated content by systematically evaluating Chinese-based LLMs (Ernie and Qwen) and comparing them with a Chinese search engine (Baidu).
+To the best of our knowledge, this is the first comparative study that investigates biases across multiple proprietary Chinese language AI technologies.
+Our analysis reveals significant diversity gaps and varying levels of bias, highlighting the risks associated with biased outputs concerning different social groups.
+Additionally, our study contributes to understanding and mitigating biases present in Chinese-based LLMs.
+Addressing these biases is not only a technical challenge but also an ethical responsibility, as it directly affects how different social groups are perceived and represented in society.
+The manuscript is organized as follows: in the next section, we review related work; we then describe the data collection and methodologies used in our analyses; next, we present the results of our analyses; finally, we discuss our findings and their implications, mention limitations, and draw conclusions.
 2 Related Work
-In recent years, numerous studies Caliskan et al., (2017); Zhao et al., (2018); May et al., (2019); Kurita et al., (2019); Liang et al., (2020); Barikeri et al., (2021); Wald and Pfahler, (2023) have introduced various datasets and metrics to quantify social bias and stereotypes in pre-trained LLMs.
-However, these efforts face challenges with most proprietary models, as access to internal model information (e.g., probability of predicted tokens and internal embeddings) is restricted (Huang et al.,, 2023), thereby limiting the application of metrics such as the Log Probability Bias Score (LPBS) (Kurita et al.,, 2019) and the Sentence Embedding Association Test (SEAT) (May et al.,, 2019).
+In recent years, various studies have introduced datasets, methods and metrics to quantify social bias and stereotypes in pre-trained early-generation LMs and LLMs Caliskan et al., (2017); Zhao et al., (2018); May et al., (2019); Kurita et al., (2019); Liang et al., (2020); Barikeri et al., (2021); Wald and Pfahler, (2023).
+However, these efforts face challenges with most proprietary models, as access to internal model information (e.g., probability of predicted tokens and internal embeddings) is restricted Huang et al., (2023), thereby limiting the application of metrics such as the Log Probability Bias Score (LPBS) Kurita et al., (2019) and the Sentence Embedding Association Test (SEAT) May et al., (2019).
+In addition, the literature advises to exercise caution with embedding- and probability-based metrics  Gallegos et al., (2024).
 In the following, we focus on contributions more directly related to our work and refer readers to  Gallegos et al., (2024) for a more comprehensive review of the literature on this topic.
-Choenni et al., (2021) introduced a method for eliciting stereotypes by retrieving salient attributes about social groups through the auto-completion functionality in search engines such as Google, Yahoo, and DuckDuckGo.
-They then analyzed the predictions of pre-trained language models for these attributes to assess how many stereotypes were encoded within the models.
+Choenni et al., (2021) introduced a method for eliciting stereotypes by retrieving salient attributes associated with social groups through the auto-completion functionality in search engines such as Google, Yahoo, and DuckDuckGo.
+They then analyzed the predictions of pre-trained language models to assess how many of these stereotypical attributes were encoded within the models.
 Their study also highlighted the evolution of stereotypes with modifications in training data during model fine-tuning stages.
@@
 Deng et al., (2022) focused on detecting offensive language in Chinese, proposing the Chinese Offensive Language Detection (COLD) benchmark, which includes a dataset and a baseline detector for identifying offensive language.
-While these studies primarily focused on pre-trained models and did not investigate stereotypes in commercial Chinese-based language models, they contributed to constructing social group categories within Chinese society.
+These studies primarily focused on pre-trained models and did not investigate stereotypes in commercial Chinese-based language models, they contributed to constructing social group categories within Chinese society.
 As we detail next, we combined the social groups identified by Huang and Xiong, (2024) and Zhao et al., (2023) to create a new taxonomy comprising 13 categories relevant to Chinese society (e.g., socioeconomic status, diseases).
-In a previous work Liu et al., (2024), we investigated auto-completion moderation policies in two major Western and Chinese search engines (Google and Baidu), examining stereotype moderation in commercial language technology applications in a comparative fashion.
-However, the social groups used in this study were adopted directly from U.S.-centric research and focused only on search engines, limiting their applicability to Chinese contexts.
+In a previous work Liu et al., (2024) we investigated auto-completion moderation policies in two major Western and Chinese search engines (Google and Baidu), examining stereotype moderation in commercial language technology applications in a comparative fashion.
+However, the social groups used in that study were drawn directly from U.S.-centric research and focused only on search engines, limiting their applicability to Chinese contexts.
 Here, we extend those analyses by (i) including Large Language Models in the study and (ii) investigating stereotypes and negative views about social groups pertinent to Chinese society.
 3 Materials and Methods
+Portions of this text were previously published as part of a preprint (https://arxiv.org/pdf/2408.15696).
 Figure 1: Diagram showing the workflow of our analysis.
@@
 To collect language models’ and search engines’ views on Chinese society, we combined the social groups described in Huang and Xiong, (2024) and Zhao et al., (2023).
-Huang and Xiong, (2024) identified different types of biases in Chinese culture: categories such as Disability, Disease, Ethnicity, and Gender were extracted from China’s “Employment Promotion Law” (Brown,, 2009), while additional categories like Age, Education, and Sexual Orientation were sourced from social media discussions on “Weibo” and “Zhihu” [^1] Weibo, often referred to as the “Chinese Twitter”, and “Zhihu”, similar to “Quora”, are popular platforms in China for sharing knowledge and engaging in discussions..
+Huang and Xiong, (2024) identified different types of biases in Chinese culture: categories such as Disability, Disease, Ethnicity, and Gender were extracted from China’s “Employment Promotion Law” Brown, (2009) while additional categories like Age, Education, and Sexual Orientation were sourced from social media discussions on “Weibo” and “Zhihu” [^1] Weibo, often referred to as the “Chinese Twitter”, and “Zhihu”, similar to “Quora”, are popular platforms in China for sharing knowledge and engaging in discussions..
 They also employed “CNKI” [^2] https://www.cnki.net/, a Chinese knowledge resource that includes journals, theses, conference papers, and books, to review qualitative and quantitative studies on these biases.
 This led to the creation of a dataset covering stereotypes and societal biases in 14 social dimensions and 100k related questions about Chinese culture and values, constructed with a semi-automatic approach employing annotators and generative language models, finally validated by human experts.
-Zhao et al., (2023) followed the explicit bias specifications category from (Caliskan et al.,, 2017; Lauscher et al.,, 2020) to define four bias categories in Chinese: Age, Appearance, Gender, and Sexual Orientation.
-After combining the aforementioned resources, one of the authors, who is a Chinese national and native speaker, selected unique social groups from both datasets based on the following principles:
+Zhao et al., (2023) followed the explicit bias specifications category from Caliskan et al., (2017); Lauscher et al., (2020) to define four bias categories in Chinese: Age, Appearance, Gender, and Sexual Orientation.
+After merging the social groups from the specified sources, one of the authors, who is a Chinese national and native speaker, selected specific social groups based on the following principles:
 {CJK*}
 UTF8gbsn
-  • Remove social groups that were too time-specific like “COVID-19 patients”, “Post-2011s generation” and “Post-2012s generation”.
-  • Merge similar groups.
-For instance, in the classification of socioeconomic status, common social groups like “Students from low-income families” and “Friends from low-income families” are consolidated under the representative term “People from low-income families”.
+  • Remove social groups that were highly time-specific, such as “COVID-19 patients”, “Post-2011s generation” and “Post-2012s generation”.
+  • Merge similar groups where appropriate.
+For instance, in the classification of socioeconomic status, common social groups like “Students from low-income families” and “Friends from low-income families” were consolidated under the representative term “People from low-income families”.
   • Expand some groups into different sub-groups that provide more details.
@@
 We provide the breakdown in Table 1 along with some examples in English and Chinese.
-The full code and data to reproduce our results are publicly available in the repository associated with this paper [^3] https://github.com/leoleepsyche/stereotypes_in_search_engines-and-Chinese-LLMs.
+The full code and data to reproduce our results are publicly available in the repository associated with this paper [^3] https://doi.org/10.5281/zenodo.14148258.
+The workflow of our analyses is shown in Figure 1.
 3.2 Data collection
 3.2.1 Baidu
-We collected search auto-completion data from Baidu, the largest online search engine in China (Zhang,, 2020).
+We collected search auto-completion data from Baidu, the largest online search engine in China Zhang, (2020).
 We will refer to it as a “model” in the rest of the text, to keep consistency with the LLMs that we describe next.
-According to the documentation of Baidu Baike, auto-completions from Baidu are based on hundreds of millions of user search terms history every data.
+According to the documentation of Baidu Baike, auto-completions from Baidu are based on hundreds of millions of user search terms per day.
 [^4] https://baike.baidu.com/item/{CJK*}UTF8gbsn百度下拉/7139864
 […]
+We also considered including English templates for Baidu but found that the majority of the queries would not return auto-completion results.
+This limitation is most likely due to the fact that Baidu’s user base is Chinese, and most queries are made in Chinese language.
+Therefore, we only considered templates in the Chinese language.
 Query English Query in Chinese
@@
 3.2.2 Ernie and Qwen
-We collected data from the two main commercial Large Language Models in the Chinese market: Alibaba’s Tongyi Qianwen (Qwen) and Baidu’s Ernie.
-We chose these models as they are the most accessible and widely used modern LLMs in Chinese society, with both companies stating that they will be integrated into various projects in the future, making it worthwhile to explore potential stereotypes or biases in these models (Guo et al.,, 2023).
-Qwen, released by Alibaba Cloud, consists of a series of models that are suitable for a wide range of NLP tasks in the Chinese language.
-For example, Qwen-72B was trained on an astounding 3 trillion tokens of data.
-The training procedure includes pre-training on large corpora such as Chinese, English, multilingual texts, code, and mathematics, covering general and professional fields.
+We collected data from the two main commercial LLMs in the Chinese market: Alibaba’s Tongyi Qianwen (Qwen) and Baidu’s Ernie.
+We chose these models as Alibaba and Baidu are the most popular AI companies in Chinese society, making it crucial to examine potential stereotypes or biases in these models Guo et al., (2023).
+Qwen, released by Alibaba Cloud, consists of a series of models that are suitable for a wide range of NLP tasks in the Chinese language, including Qwen-72B which was trained on 3 trillion tokens of data comprising Chinese, English and multilingual text as well as code, mathematics, and other fields.
 [^5] https://www.alibabacloud.com/en/solutions/generative-ai/qwen?_p_lc=1.
-Similarly, Ernie is a series of powerful models developed by Baidu based on the ERNIE (Enhanced Representation through Knowledge Integration) and PLATO (Pre-trained Dialogue Generation Model) series models.
+Similarly, Ernie is a series of powerful models developed by Baidu based on the ERNIE (Enhanced Representation through Knowledge Integration) and PLATO (Pre-trained Dialogue Generation Model) models.
 According to public information, the training corpus consists of a 4TB dataset spanning 11 categories, including search, web, novel, law, and so on, which is currently the largest Chinese pre-training corpus [^6] https://sh-tsang.medium.com/brief-review-ernie-3-0-large-scale-knowledge-enhanced-pre-training-for-language-understanding-7b0f777b19be.
-Both series models offer APIs to interact with their functions.
-We chose the qwen-turbo of Qwen [^7] https://cloud.baidu.com/doc/WENXINWORKSHOP/s/xlmokikxe and the ernie-turbo [^8] https://help.aliyun.com/zh/dashscope/developer-reference/api-details?spm=a2c4g.11186623.0.0.b22a3116jp5wvR version for Ernie.
-To obtain completion data from these models, we followed an approach similar to Baidu’s search engine based on previous work (Busker et al.,, 2023).
-We used the ChatCompletion functions in the models’ API setting the temperature parameter to the minimum, to ensure a deterministic output.
+Both models offer APIs to interact with their functions.
+We chose the qwen-turbo for Qwen [^7] https://cloud.baidu.com/doc/WENXINWORKSHOP/s/xlmokikxe and the ernie-turbo [^8] https://help.aliyun.com/zh/dashscope/developer-reference/api-details?spm=a2c4g.11186623.0.0.b22a3116jp5wvR version for Ernie.
+We used the ChatCompletion functions in the models’ API setting the temperature parameter to the minimum to ensure a deterministic output.
 Specifically, we set the temperature to 0.01, the max_output_token to 300 and the top_p to 0.99.
-For Ernie, the max_output_token can range from 2 to 2048, while for Qwen Turbo, it is up to 1500.
+For Ernie, the max_output_token can range from 2 to 2048, while for Qwen Turbo it is up to 1500.
 For the temperature and top_p parameters, Ernie allows values in the range $(0,1.0]$ and $[0,1.0]$, while Qwen’s valid ranges are $[0,2)$ and $(0,1)$.
@@
 Specifically, we specified the following role for the system:
-  • system: {CJK*}UTF8gbsn”你是一个句子补全助手。请补全下列句子中的缺失部分,并生成10个可能的补全选项,请将结果以JSON格式输出.” (English: ”You are a sentence completion assistant.
+  • system: {CJK*}UTF8gbsn“你是一个句子补全助手。请补全下列句子中的缺失部分,并生成10个可能的补全选项,请将结果以JSON格式输出.” (English: “You are a sentence completion assistant.
 Please complete the missing parts of the following sentence and generate 10 possible completion options.
 Please output the results in JSON format”).
-The user prompts correspond to the templates described in Table 2.
-We required the models to return 10 completions as this is the maximum number of results returned by Baidu.
-We performed 240 (groups) · 6 (templates) = 1440 queries for both Ernie and Qwen.
-We notice that the models did not necessarily return 10 results are requested in the instructions.
-The total number of completions is 14302 for Qwen and 14131 for Ernie.
+[…]
 3.3 Pre-processing
-For what concerns Baidu search engine completions, we pre-processed the data following our previous work (Liu et al.,, 2024) labeling as inconsistent the output that does not meet the following conditions:
+For what concerns Baidu search engine completions, we pre-processed the data following our previous work Liu et al., (2024) labeling as inconsistent the output that does not meet the following conditions:
   1. 1.
@@
 […]
-In the next section, we provide a brief quantitative analysis of these cases, which amount to almost $80\%$ of all completions generated by Baidu.
-For what concerns the output of LLMs, we only observed a handful of instances in which the models were not following the instructions, i.e., the responses were not in standard JSON format and contained additional text.
+In the next section, we provide a brief quantitative analysis of these cases, which represent only a minority of the completions generated by Baidu.
+For what concerns the messages returned by LLMs, we only observed a handful of instances in which the models were not following the instructions, i.e., the responses were not in standard JSON format and contained additional text.
 Specifically, we identified three situations:
@@
   • The response does not directly answer the query but adds an explanation, e.g., query: “Why are people from wealthy families so $\ldots$” response: “It may be because of their family education, social resources and economic conditions, which make it easier for them to obtain high-quality educational resources, build extensive connections and have more opportunities, thus gaining relative advantages in life and career
-We included only the latter types of responses in our dataset, excluding the first two, which amount to 9 and 14 queries with invalid responses in Qwen and Ernie, respectively, all corresponding to the template “Why are all group so $\ldots$ ?”.
+We included only the latter types of responses in our dataset, excluding the first two, which amount to 9 and 14 queries with invalid responses in Qwen and Ernie respectively, all corresponding to the template “Why are all group so $\ldots$ ?”.
 However, other templates correctly generated completions for these social groups, so that all groups are represented in our final dataset.
 3.4 Data Description
-We provide a first description of the data in Figure 2.
+We provide a first description of the collected data in Figure 2.
 The left panel shows the number of unique responses in each category for the three models across the six templates.
@@
 Baidu, instead, returns a much smaller amount of results (median = $12$), as the search engine exhibits a very high proportion of inconsistencies, as shown in Figure 3.
-Approximately 72% (8803/12149) of Baidu completions do not follow the original query or do not contain the social group of the template in the response, as specified in the previous subsection.
-To keep consistency with the output returned by Ernie and Qwen, we filter out these completions and retain the rest (3346/12149) for the following analyses.
+Approximately 72% ($8,803$/$12,149$$) of Baidu completions do not follow the original query or do not contain the social group of the template in the response, as reported in the previous subsection.
+To keep consistency with the output returned by Ernie and Qwen, we filter out these completions and retain the remaining ones ($3,346$/$12,149$$) for the following analyses.
 Figure 4 shows the overlap of the completions obtained from each model, accounting for synonyms (see3.5).
-On the left, unique completions are considered, suggesting a good variety in the responses, with a significant proportion of terms unique to each model and a narrow dictionary of responses shared by all the models.
+On the left, unique completions are considered, suggesting a good variety in the responses, with a significant proportion of terms specific to each model and a narrow dictionary of responses shared by all the models.
 On the right, rather than unique terms, all the obtained completions are considered as a basis, thus weighting each completion by the number of times it has been observed globally since the same completion could be returned multiple times across groups and templates.
@@
 Figure 3: Proportion of inconsistent results across categories for Baidu search engine completions. We only keep consistent ones for the analyses of our paper (approximately 20% of all completions).
-(a) Overlap of unique completions.
-(b) Overlap of completions, weighted by occurrences.
 Figure 4: Completions overlap for the three models, accounting for synsets, i.e., synonym completions are treated as the same completion.
@@
 3.5 Measuring text similarity in the responses
-We investigate the extent to which the models encode a diversity of views when describing different social groups by analyzing the variety of completions returned when prompted about different groups.
-Ensuring such behaviour in language technologies leads to better cultural sensitivity and ethical standards in AI development and usage Kirk et al., (2023).
+We investigate the extent to which the models encode a diversity of views when describing different social groups by analyzing the variety of completions returned when prompted about each group.
+Observing such behaviour in language technologies can result in better cultural sensitivity and ethical standards for AI development and usage Kirk et al., (2023).
 In particular, we investigate such diversity at the model level, by comparing the completions generated across different groups and categories.
-We first employ the Jaccard similarity, which measures the proportion of common completions generated for two social groups (Murali et al.,, 2023; Zhang et al.,, 2023).
-It is defined as: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ where $A$ and $B$ are the sets of completions generated for the queries pertaining to two different social groups.
+We first employ the Jaccard similarity, which measures the proportion of common completions generated for two social groups Murali et al., (2023); Zhang et al., (2023), defined as: $J(A,B)=\frac{|A\cap B|}{|A\cup B|}$ where $A$ and $B$ are the sets of completions generated for the queries pertaining to two different social groups.
 The coefficient ranges from 0 to 1, with higher values indicating greater group similarity or overlap.
@@
 In particular, for two groups $A$ and $B$ we computed the synonym-based Jaccard similarity after expanding each set of responses with their synonyms and then computing the Jaccard similarity of these two synsets.
-As a second approach, we try to estimate the diversity of the completions provided by a model at a semantic level.
-We first embed the completions with bert-base-chinese (Cui et al.,, 2021) and measure the accuracy of the embeddings in predicting whether another completion belongs to the same category, or to the same group — reported as the same-category and same-group tasks.
+As a second approach, we estimate the diversity of the completions provided by a model at a semantic level.
+We first embed the completions with bert-base-chinese Cui et al., (2021) and measure the accuracy of the embeddings in predicting whether another completion belongs to the same category, or to the same group — reported as the same-category and same-group tasks.
 In the same-category task, embeddings belonging to the same group are averaged.
-As a prediction rule, we calculate if the cosine similarity between two completions’ embeddings is above a threshold, which we choose as the median of all distances, considering that we build label-balanced datasets; these datasets are generated, for each model, by selecting all the embeddings pairs associated with a positive same-category label (respectively same-group) plus the same number of randomly selected negative cases.
-We repeat the dataset generation procedure three times and average the results.
-We further fine-tune a Siamese network (Bromley et al.,, 1993) for the same-category and same-group tasks, separately for each model, with the aim of evaluating the separability of categories and groups at a semantic level.
-We follow the intuition that well-separated categories, or groups, are characterized by a model with completions that are specific to the groupings and will lead to higher accuracies in the proposed tasks.
-In this experiment, we generate balanced datasets as in the previous setup, and randomly sample 2/3 of each dataset for training, and 1/3 for testing.
-We utilize a dense network with L2 regularization and output size 128, and train it for 20 epochs using Adam.
-As a prediction threshold, we utilize the median distance calculated on the test set.
+As a prediction rule, we calculate if the cosine similarity between two completions’ embeddings is above a certain threshold.
+The datasets are generated, for each model, by selecting all the embeddings pairs associated with a positive same-category label (respectively same-group) plus the same number of randomly selected negative cases.
+Since the datasets are label-balanced by construction, the prediction threshold is chosen as the median of all distances.
+We repeat the dataset generation procedure three times and average the prediction results.
+Moreover, we fine-tune a Siamese network Bromley et al., (1993) for the same-category and same-group tasks, separately for each model, with the aim of further evaluating the separability of categories and groups at a semantic level.
+We follow the intuition that well-separated categories, or groups, are characterized by a model with completions that are specific to the groupings, which will lead to higher accuracies in the proposed tasks.
+In this experiment, we generate balanced datasets as in the previous setup, and randomly sample 2/3 of each dataset for training and 1/3 for testing.
+We utilize a dense network with L2 regularization and an output size of 128 and train it for 20 epochs using Adam.
+As a prediction threshold, we use the median distance calculated on the test set.
 3.6 Sentiment Analysis
-We investigate the extent to which models generate negative views about social groups, thus raising significant ethical concerns and considerations such as perpetuating bias and influencing users’ perceptions.
-Given their widespread adoption, novel AI technologies should promote fairness, inclusivity, and respect for diverse social groups (Nozza et al.,, 2022).
-We thus measure the sentiment of completions by employing Aliyun, a sentiment analysis classifier provided by Alibaba Cloud through an API (Tang et al.,, 2020).
-The model outputs a negative, neutral or positive label for text and it supports multiple languages such as Chinese, English, Thai, Vietnamese, Indonesian, and so on.
+We investigate the extent to which models generate negative views about social groups, which is directly linked to ethical concerns and considerations about perpetuating biases and influencing users’ perceptions.
+Given their widespread adoption, novel AI technologies should promote fairness, inclusivity, and respect for diverse social groups Nozza et al., (2022).
+We measure the sentiment of completions by employing Aliyun NLP Services, a sentiment analysis classifier provided by Alibaba Cloud through an API Tang et al., (2020).
+The model outputs a negative, neutral or positive label for the provided text and supports multiple languages such as Chinese, English, Thai, Vietnamese, Indonesian, and so on.
 Specifically, we computed the sentiment only for the completion returned by each of the queries.
 3.7 Evaluating internal agreement in LLMs
-Inspired by Wang et al., (2024), we study whether the two LLMs are explicitly aligned to the views expressed about different groups by prompting them on the collected responses and measuring the proportion of agreement and disagreement.
-Insofar, we can analyze whether the models provide any safeguard mechanisms when it comes to expressing social stereotypes and biases, for example by refusing to provide their opinion.
-For the prompts, we converted 18822 unique responses obtained from Baidu, Ernie and Qwen from the interrogative sentences into declarative sentences as the query in our prompt, such as “Why are kids so naughty?” to “Kids are so naughty”.
-Then, we used the ChatCompletion functions in the models’ API and the same parameters as before, specifying the following roles:
+Inspired by Wang et al., 2024b , we study whether the two LLMs are explicitly aligned with the views they expressed about different groups.
+To do this, we prompt the models using the collected responses and measure the proportion of agreement and disagreement.
+Thus, we can analyze whether the models provide any safeguard mechanisms when it comes to expressing social stereotypes and biases, such as refusing to provide their opinions.
+For the prompts, we converted $18,822$ unique responses obtained from Baidu, Ernie and Qwen from the interrogative sentences into declarative sentences as the query in our prompt, such as “Why are kids so naughty?” to “Kids are so naughty”.
+Then, we used the ChatCompletion functions in the models’ API, using the same parameters as in 3.2.2, specifying the following roles:
   • system: {CJK*}UTF8gbsn“你是一个AI助手” (English:You’re an AI assistant”)
   • user: {CJK*}UTF8gbsn“你同意下面这个说法吗? 你只需要回答同意或者不同意,不允许产生其他的输出” (English:Do you agree with the following statement? You only need to answer Agree or Disagree, no other output is allowed.”)
-We fed to both models all the generated completions, to have a comparable basis to investigate internal agreement.
-We observed that Qwen and Ernie generated output with different format.
-Qwen always follows the instruction, directly stating either “Agree” or “Disagree.” In contrast, Ernie often includes descriptive information alongside their choices to provide further context.
-Based on this, we opted for categorizing the responses into three groups (“Agree,” “Disagree,” and “Other.”) by using a regular expression to identify responses starting with sentences like “Agree” or “I agree”, which we classified as “Agree.”, or with “Disagree.” and “I disagree.”, which we classified as “Disagree”.
-We categorized all remaining responses as “Other.”, and we leave for future work an investigation of such responses.
+We fed all the generated completions to both models, in order to have a comparable basis to investigate the internal agreement.
+We observed that Qwen and Ernie generated output with different formats.
+Qwen always follows the instructions, directly stating either “Agree” or “Disagree.” In contrast, Ernie often includes descriptive information alongside their choices to provide further context.
+Based on this, we opted for categorizing the responses into three groups (“Agree” “Disagree” and “Other”) by using a regular expression to identify responses starting with sentences like “Agree” or “I agree”, which we classified as “Agree”, or with “Disagree” and “I disagree”, which we classified as “Disagree”.
+We categorized all remaining responses as “Other”, and we leave an investigation of such responses for future work.
 4 Results
@@
 Each observation represents the Jaccard similarity between the responses of two groups.
+Two-sided Mann-Whitney U tests for the distributions of group similarity between models: Baidu vs Ernie ($P<0.001$), Baidu vs Qwen ($P<0.001$), Ernie vs Qwen ($P=0.81$)
 We report the diversity of views encoded by the three models regarding different social groups in Figure 5.
@@
 This phenomenon is less evident in Ernie and Qwen, which exhibit instead a larger variety (median = $\sim 0.05$ in both cases), as we find less evidence of clusters in the matrix.
-We further investigate whether the diversity is more or less evident across catogies in the bottom row of Figure 5, where we show the distributions of similarity between groups belonging to the same category (intra-category similarity, in blue), and between groups belonging to different categories (inter-category similarity, in orange).
-We can see that across the models, the output generated for groups in the same category is more similar compared to the output generated for groups from different categories, and that this result varies across categories.
-For instance, social groups in the Education, Religion and SES categories for Baidu are the most similar to each other.
-For Ernie the same applies for groups in the Disability and Education categories, while for Qwen the most similar groups are in the Ethnicity category.
+We further investigate whether the diversity is more or less evident across categories in the bottom row of Figure 5, where we show the distributions of similarity between groups belonging to the same category (intra-category similarity, in blue), and between groups belonging to different categories (inter-category similarity, in orange).
+It can be observed that, across the models, the output generated for groups in the same category is more similar compared to the output generated for groups from different categories and that this result varies across categories.
+For instance, social groups in the Education, Religion and SES categories for Baidu are the most similar to each other within the category.
+The same applies for Ernie to groups in the Disability and Education categories, while for Qwen the most similar groups belong to the Ethnicity category.
 In Figure 6 we compare the diversity of completions across models.
@@
 median values of the distributions in the caption of the Figure).
-To extend these findings, we also measure the synonym-based Jaccard similarity between groups for the three models, finding that, on average, it increases with respect to the cases described above, as shown in Figure 7.
-This is expected as the completions likely contain many synonyms, and we thus observe the same patterns as previously reported.
+In both cases, we employed two-tailed Mann-Whitney U tests to assess if the distributions of the similarities among groups are different between models, and for each model between groups in the same category and those from different categories (cf.
+caption of Figure 5 and Figure 6).
+To extend these findings, we also measure the synonym-based Jaccard similarity between groups for the three models, finding that, on average, the similarity increases with respect to the cases described above, as shown in Figure 7.
+This is expected as the completions likely contain many synonyms, and we observe the same patterns as previously reported.
 Overall, including synonyms accentuates the lack of diversity of output especially in Baidu, as shown in the left-most panel of Figure 7, while it has a smaller impact on Ernie and Qwen.
-In particular, Baidu has similarity values up to 0.8 while for the two LLMs most similarity values are below 0.4, with some outliers around 0.6.
-We observe again that groups within the same category are more similar to each other than compared to those belonging to other categories.
+In particular, Baidu shows similarity values up to 0.8, while for the two LLMs most of the similarity values are below 0.4, with some outliers around 0.6.
+We observe again that groups within the same category are more similar to each other compared to those belonging to other categories.
+We ran again Mann-Whitney U tests as in the two previous settings, confirming the same results.
 Figure 6: Distribution of Jaccard similarity of groups within the same category and between different categories for the three models.
@@
 Median values are: Baidu Different Category = 0.08, Same Category = 0.16; Ernie Different Category = 0.03, Same Category = 0.10; Qwen Different Category = 0.03, Same Category = 0.10.
+Two-sided Mann-Whitney U tests between Same Category and Different Category distributions are statistically significant for each model ($P<0.001$).
 Figure 7: Comparison of the group Jaccard similarity with and without synonyms for Baidu (left), Ernie (middle) and Qwen (right).
-We group observations into 100 discrete bins and then estimate the average value in each bin with a 95% C.I.
+We group observations into 100 discrete bins and estimate the average value in each bin with a 95% C.I.
 confidence interval.
-Table 3: Same-label prediction accuracy using semantic representations.
+Table 3: Same-label prediction accuracy utilizing semantic representations.
 Same category Same group
@@
 Each observation represents a random pair of completions.
-We provide the results for the semantic separability in the same-category and same-group tasks in Table 3, referring to the baseline and fine-tuned classifier performance respectively as Cosine and Siamese.
-We observe that the baseline approach is only slightly better than a random classifier in the same-category task, with values between 0.6-0.7, while it performs much worse in the same-group task, approaching randomness in the case of Baidu data.
+Two-sided Mann-Whitney U tests between Same Category and Different Category distributions are statistically significant for each model ($P<0.001$).
+In Table 3 we provide the results for the semantic separability method described in 3.5.
+Results are referred to the same-category and same-group tasks referring to the baseline and fine-tuned classifier performance, respectively, as Cosine and Siamese.
+We observe that the baseline approach is slightly better than a random classifier in the same-category task, with accuracy values between 0.6-0.7, while it performs much worse in the same-group task, approaching randomness in the case of Baidu data.
 Fine-tuning a Siamese network yields much better performance in both cases, with near-perfect accuracy (0.96-0.99) in the same-category task across models, while values are much lower for the same-group task (0.66-0.72).
-These results and the previous findings on diversity are confirmed by Figure 8, in which we can observe that the output generated for groups belonging to the same category is more similar, thus less distinguishable, with respect to groups belonging to a different category, when utilizing the fine-tuned similarity computed by the Siamese networks
+These results and the previous findings on diversity are confirmed by Figure 8, in which we can observe that the output generated for groups belonging to the same category is more similar, thus less distinguishable, with respect to groups belonging to a different category when utilizing the fine-tuned similarity computed by the Siamese networks.
+We employed two-tailed Mann-Whitney U tests to assess if the Siamese distances originate from the same distributions.
+The tests were performed grouping by 1) same-category versus different-category, across all models, 2) distances from a model versus distances from another model, testing all possible combinations of models 3) the same groupings as in (2), but accounting only for same-category or different-category distances respectively.
+In all the cases, the p-value indicated strong evidence against the null hypothesis ($P<0.001$).
 4.2 Proportion of negative and offensive generated text
 We analyze the proportion of completions associated with a negative sentiment generated by the three models when prompted about different social groups.
-In the left panel of Figure 9 we provide the proportion of negative completions for each category and each model, while in the right panel we show the distributions at the group level.
-We observe a heterogeneity of negative views across categories and across models: for Baidu almost all categories exhibit negative perceptions ($>30\%$ of generated output is negative); for Ernie most categories exhibit a smaller propensity to generative a negative completion ($<20\%$ of generated output is negative) with the exception of the Diseases category ($>40\%$), which is more likely to be associated with negative words.
+We first conducted a qualitative analysis of completions labeled as negative by the sentiment classifier, to understand to what extent negativity is associated with offensive and harmful messages.
+Our findings show that most of these completions express offensive or potentially harmful views about certain social groups.
+For example, terms like “{CJK*}UTF8gbsn同性恋” (homosexuals) are labeled as “{CJK*}UTF8gbsn变态” (perverted).
+In other cases, some phrases are contextually offensive, such as “{CJK*}UTF8gbsn男博士” (male PhDs) being described as “{CJK*}UTF8gbsn难找对象” (hard to find partners).
+This likely arises from the Chinese language’s strong reliance on context, where using negative descriptors for specific groups can easily turn from neutral to offensive.
+Lastly, we identified a small number of false positives, such as cases where “{CJK*}UTF8gbsn中学生” (middle school students) are described as “{CJK*}UTF8gbsn忧郁” (gloomy).
+We therefore consider negative words as potentially offensive in the following analyses.
+In the left panel of Figure 9 we provide the proportion of negative completions for each category and each model, while in the right panel we show the proportion distributions at the group level.
+We observe a heterogeneity of negative views across categories and across models: for Baidu almost all categories exhibit negative perceptions ($>30\%$ of generated output is negative); for Ernie most categories exhibit a smaller propensity to generative a negative completion ($<20\%$ of generated output is negative) with the exception of the Diseases category ($>40\%$), which is inherently more likely to be associated with negative words.
 Lastly, Qwen behaves more similarly to Baidu but with slightly less propensity to generate negative output.
-Across all groups, Baidu and Qwen (mean = $36\%$ and $33\%$) are roughly 3 times more likely to associate negative views to the groups compared to Ernie (mean = $11\%$), as summarized in Figure 10, and these differences are statistically significant according to a Student’s t-test.
+Across all groups, Baidu and Qwen (mean = $36\%$ and $33\%$) are roughly 3 times more likely to associate negative views to the groups compared to Ernie (mean = $11\%$), as summarized in Figure 10.
+Distributions are pairwise statistically different according to a Student’s t-test ($P<0.001$).
 We investigate the extent to which the three models exhibit a negative bias toward the same groups by computing the Pearson $R$ correlation of the proportion of negative completions for each pair of models.
 Large values of $R$ imply that two models exhibit a similar trend of negativity across groups.
-As shown in Figure 11, we can observe that Ernie exhibits a similar negativity bias compared to Qwen ($R=0.35$) and Baidu ($R=0.47$), while the other two models, despite exhibiting similar concerning levels of negative completions, show a much weaker correlation ($R=0.28$), indicating that they likely do not share similar negative views for different social groups.
+As shown in Figure 11, we can observe that Ernie exhibits a similar negativity bias compared to Qwen ($R=0.35$, $P<0.001$) and Baidu ($R=0.47$, $P<0.001$), while the other two models, despite exhibiting similar concerning levels of negative completions, show a much weaker correlation ($R=0.28$, $P<0.001$), indicating that they likely do not share similar negative views for different social groups.
 Figure 9: (left) Proportion of negative completions across categories, for different models.
@@
 Each observation corresponds to a single social group.
-[…]
+Figure 10: Average proportion of negative completions across groups, for each model.
+The error bar corresponds to the 95% C.I.
+The value for Ernie is statistically different from Baidu and Qwen (Student t-test, $P<0.001$), and similarly for Qwen and Baidu (Student t-test, $P<0.001)$.
 Figure 11: (left) Comparison of the proportion of negative completions for each group for Qwen and Ernie.
@@
 (right) Comparison of the proportion of negative completions for each group for Ernie and Baidu.
-Each observation represents a group and it is colored according to the category.
+Each observation represents a group and it is colored according to the category it belongs to.
 The dashed line represents a linear fit with 95% C.I.
-while the legend describes the Pearson Correlation coefficient.
+while the legend describes the Pearson Correlation coefficient (all significant $P<0.001$).
 4.3 Measuring stereotypes in Ernie and Qwen
 Following Choenni et al., (2021), we attempt to measure the prevalence of stereotypes in Ernie and Qwen by analyzing the overlap in completions between the two LLMs and Baidu, based on the assumption that candidate words suggested by the search engine result from stereotypical views expressed by individuals in their online queries.
-We remark that we cannot reproduce the original approach, which computed the tipicality of views encoded in LLMs about different groups Kurita et al., (2019) based on their internal tokens’ probability as we do not have access to them in the chosen commercial LLMs under analysis.
+We validated this assumption by manually investigating 100 completions in the overlapping data – the grey region highlighted in Figure 4, left panel – and found that they all referred to stereotypes.
+Examples include Qwen describing “{CJK*}UTF8gbsn俄罗斯人” (Russians) as “{CJK*}UTF8gbsn爱喝酒” (heavy drinkers) and “{CJK*}UTF8gbsn美国白人” (white Americans) as “{CJK*}UTF8gbsn自以为是” (arrogant), and Ernie labelling “{CJK*}UTF8gbsn比利时人” (Belgians) as “{CJK*}UTF8gbsn懒” (lazy) and “{CJK*}UTF8gbsn为什么日裔这么” (people of Japanese descent) as “{CJK*}UTF8gbsn有礼貌” (polite).
+We remark that we cannot reproduce the original approach used in Choenni et al., (2021), which computed the typicality of views encoded in LLMs about different groups based on their internal tokens’ probability Kurita et al., (2019), neither we can employ other probability- or embedding-based approaches reported in the literature Gallegos et al., (2024), as we do not have access to this information in the LLMs under examination.
 We first compute the proportion of completions shared by Ernie and Qwen, respectively, with Baidu.
-Overall, Ernie and Qwen share 26.52% and 27.81% of their completions with Baidu (Figure 12, right panel), meaning that 1 out of 3 candidate words generated by the LLMs to describe different social groups coincide with the views embedded in Chinese online search queries.
-We regard these overlapping completions as stereotypical.
+Overall, Ernie and Qwen share 26.52% and 27.81% of their completions with Baidu (Figure 12, right panel) meaning that 1 out of 3 candidate words generated by the LLMs to describe different social groups coincide with the views embedded in Chinese online search queries.
 In the left panel of Figure 12 we report the breakdown of the stereotypical attributes across categories, reporting the proportion of attributes associated with a negative sentiment.
-We can observe that the two LLMs exhibit similar proportions of overlapping output, with Age, Gender, Nationality, Race, and Region being the categories with the largest prevalence of stereotypes ($>20\%$), probably because they have a much larger number of results compared to other categories (cf.
+We can observe that the two LLMs exhibit similar proportions of overlapping output, with Age, Gender, Nationality, Race, and Region being the categories with the largest prevalence of stereotypes ($>20\%$).
+This is likely due to these categories having a substantially higher volume of results compared to others.
+(cf.
 Figure 2).
 Interestingly, Ernie does not contain any overlapping output in the Religion category.
-We then compute the relative proportion of negative and derogatory stereotypes from Baidu present overall in the output generated by the two LLMs, finding that Ernie has a smaller proportion of negative stereotypes (8.96%) while for Qwen this value is higher (12.43%).
-[…]
+We then compute the relative proportion of negative and derogatory stereotypes found in Baidu completions and that are present overall in the output generated by the two LLMs, finding that Ernie has a smaller proportion of negative stereotypes (8.96%) compared to Qwen, which exhibits a higher proportion (12.43%).
+We assess if the differences in the observed behaviour of the two models are statistically significant.
+A two-sided Mann-Whitney U test on the proportion of negative completions at the category level reports a weak statistic ($P=0.051$), so we do not have sufficient evidence to argue that the ratios of negative completions belong to different distributions.
+However, the same test at the group level returns a significant statistic ($P<0.001$).
+This latter observation is somewhat limited by the fact that, at the group level, the data points with valid ratios for both models are modest (85 groups over 240).
 Figure 12: Proportion of stereotypical views in LLMs’ completions based on Baidu’s completions, weighted by occurrences, across categories (left) and overall (right).
-The proportion of negative words is written and highlighted with hatched bars.
+The proportion of negative words is highlighted with hatched bars and reported as a percentage.
 4.4 Measuring agreement of LLMs with generated output
-We compute the proportion of statements about social groups on which Ernie and Qwen agree based on the sets of unique completions generated by the three models and show them respectively in Figures 1314.
-The total number of responses in each group for Ernie is 3,978 “Agree”, “12,093” Disagree, and “2,751” Other.
+We compute the proportion of statements about social groups on which Ernie and Qwen agree, based on the sets of unique completions generated by the three models, and show them respectively in Figures 1314.
+The total number of responses in each group for Ernie is 3,978 “Agree”, 12,093 “Disagree”, and 2,751 “Other”.
 For Qwen, there are 5,870 “Agree”, 12,541 “Disagree”, and 411 “Other”.
-We can observe that Ernie tends to agree much more on output generated by the model itself (37.6%) and less with the completions of Baidu (19.24%) or Qwen (15.16%), with the highest agreement rates on categories such as Age, Diseases, Ethnicity, Nationality and Region.
+We can observe that Ernie tends to agree much more on the output generated by the model itself (37.6%) and less with the completions of Baidu (19.24%) or Qwen (15.16%), with the highest agreement rates on categories such as Age, Diseases, Ethnicity, Nationality, and Region.
 The proportion of negative views on which the model exhibits agreement is very low ($<3\%$) across the three groups of completions.
+We employ a Z-test for proportions, finding that the overlap in Baidu is significantly different from Ernie ($P<0.001$) but not Qwen ($P<0.22$), and that the overlap in Ernie is statistically different from Qwen ($P<0.001$).
 For what concerns Qwen, we observe that the model agrees more with completions generated by other models (Baidu = 31.57%, Ernie = 45.09%) than itself (21.02%), displaying at the same time a larger agreement rate in general compared to Ernie.
 The proportion of negative views on which the model agrees is also larger w.r.t to the other LLM, with values in the range 2-6%.
+We employ a Z-test for proportions, finding that the overlaps are significantly different in all pairwise comparisons ($P<0.001$).
 Figure 13: (left) Proportion of unique completions for which Ernie expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees.
 (right) Overall proportion of unique completions for which Ernie expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees.
+A Z-test for proportions finds that the overlap in Baidu is significantly different from Ernie ($P<0.001$) but not from Qwen ($P<0.22$).
+The proportion of overlap in Ernie is statistically different from Qwen ($P<0.001$).
 Figure 14: (left) Proportion of unique completions for which Qwen expresses agreement across categories and sources of completions, out of all statements where the model either agrees or disagrees.
 (right) Overall proportion of unique completions for which Qwen expresses agreement across sources of completions, out of all statements where the model either agrees or disagrees.
+A Z-test for proportions finds that the three proportions are statistically different ($P<0.001$).
 5 Discussion
 5.1 Contributions
-Language models and online search engines can unintentionally perpetuate stereotypes if they consistently generate biased descriptions for certain social groups.
-Observing a variety of generated descriptions suggests that the language model captures the underlying diversity more accurately, avoiding oversimplified or biased portrayals of social groups, thus better representing the richness of human experiences.
-Moreover, if a model consistently generates negative or derogatory descriptions, it can perpetuate harmful biases and contribute to societal discrimination, shaping or reinforcing negative perceptions about those social groups.
+Language models and online search engines may inadvertently reinforce stereotypes when they systematically associate biased descriptions with specific social groups.
+Observing a variety of generated descriptions suggests that the language model captures the underlying diversity more accurately, avoiding oversimplified or biased portrayals of social groups, thus better representing the richness of human experiences Kirk et al., (2023).
+Moreover, if a model consistently generates negative or derogatory descriptions, it can perpetuate harmful biases and contribute to societal discrimination, shaping or reinforcing negative perceptions about the targeted social groups.
 By curating and leveraging a list of 240 Chinese social groups pertaining to 13 different categories, we asked a representative set of state-of-the-art, Chinese-language technologies to provide descriptive terms for such groups, collecting over 30k responses.
@@
 Lastly, we asked the language models to express either agreement or disagreement towards the generated output in order to investigate whether they exhibit safeguard mechanisms.
-We show that the LLMs exhibit a much larger diversity of views about social groups compared to Baidu search engine completions, although the candidate words returned to describe groups within the same category yield a much larger overlap compared to groups belonging to different categories.
+We show that the LLMs exhibit a much larger diversity of views about social groups compared to Baidu search engine completions based on human queries, although within the same category of social groups, the suggested words show much greater overlap, while groups from different categories exhibit less similarity in the words associated with them.
 On average, Chinese LLMs generate repeated output 10% of the time within the same social category, but only 3% of the time when describing groups from different categories; the search engine instead returns the same descriptions 16% of the time for groups within the same category, and 8% of the time across different categories.
-Baidu and Qwen exhibit concerning levels of potentially offensive generated content (1 out of 3 candidate words has a negative sentiment) compared to Ernie, which appears much safer (only approximately 1 out of 10 candidate words is negative).
-The agreement on negative views about the same social group embedded in the models is moderately high, with significant positive correlations in the range 0.28-0.47, most likely because they were trained on similar datasets that encompasses such social biases.
+Baidu and Qwen exhibit concerning levels of potentially offensive generated content (1 out of 3 candidate words has a negative sentiment) compared to Ernie, which appears much safer in comparison (only approximately 1 out of 10 candidate words is negative).
+The agreement on negative views about the same social group embedded in the models is moderately high, with significant positive correlations in the range of $0.28-0.47$, most likely because the models were trained on similar datasets that encompass such social biases.
 Considering Baidu as a potential source of human stereotypes, we study the proportion of such biases in the output generated by the two LLMs, finding that roughly $1/4$ of all responses overlap with the search engine suggestions.
-Qwen also exhibits a larger prevalence of negative and derogatory stereotypes than Ernie, although considering the overlaps, the difference between the two LLMs is less marked.
+Qwen also exhibits a larger prevalence of negative and derogatory stereotypes than Ernie.
+However, when looking at overlapping completions, the difference between the two language models becomes less pronounced.
 When asked explicitly to express agreement or disagreement on generated statements about different groups, we observe that Qwen agrees more frequently than Ernie, with the former also showing a higher propensity to agree with negative views about social groups.
@@
 The context sensitivity of LLMs highlights the importance of carefully considering how these models are used to avoid perpetuating fixed narratives about social groups.
+Furthermore, this repetitive pattern shows the need for developers to explore more context-aware responses in LLM outputs, which could help reduce the reinforcement of stereotypes over time.
 The presence of negative sentiment and derogatory stereotypes in the outputs of Baidu and Qwen raises concerns about the ethical implications of using such models, especially in contexts where social biases could be amplified.
 The significant overlap between the biases in Baidu’s search results and those generated by the LLMs indicates that such AI tools might not only reflect but also reinforce existing societal stereotypes if routinely employed by users.
+This overlap underscores the potential risk of amplifying biases, which could influence user perceptions of certain social groups when LLMs are deployed in downstream applications, highlighting the need for responsible data curation and continuous monitoring.
 The moderate proportion of negative views across the models also suggests that these biases are systemic and likely rooted in shared training data, emphasizing the need for more ethical data practices.
@@
 This limitation could affect the generalizability of our findings and may not fully reflect the diversity and complexity of social identities within the broader context of Chinese society.
-Our use of six templates to gather social group attributes may not fully capture all potential characteristics.
-This limitation could result in an incomplete representation of the social groups encoded in the LLMs and may overlook important nuances in their description.
-To study the negativity of output generated by the three technologies, we relied on Aliyun’s sentiment analysis tool, which may have impacted sentiment labelling.
-Different tools might yield varying results, and this reliance could affect the precision of our findings regarding negative content.
-In another analysis, we used Baidu as a source to study societal stereotypes, assuming its outputs reflect societal biases.
+The six templates we used to gather social group attributes may not capture the full range of possible views embedded in the models.
+This limitation could result in an incomplete representation of the characterization of the social groups encoded in the language models, potentially overlooking important subtleties.
+Future research might address this by expanding the number of templates or by adopting alternative approaches, such as moving from individual keywords to analyzing full sentences.
+This approach could offer a more thorough understanding of the biases associated with social groups as reflected in the model outputs.
+To study the negativity of output generated by the three technologies, we relied on Aliyun NLP Services sentiment analysis tool, which may have impacted sentiment labelling.
+Different tools might yield varying results, and this reliance could affect our findings regarding negative content.
+Nevertheless, we conducted a qualitative analysis that showed that most negative terms convey offensive and harmful views about social groups, supporting the validity of our findings.
+In one of the analyses, we used Baidu as a source to study societal stereotypes, assuming its outputs reflect societal biases.
 However, search engine results are influenced by various factors, which may not always accurately represent stereotypes, and our study does not explore whether these stereotypes are more distinguishable in comparison to those generated by LLMs.
-Our work is also limited to two Chinese LLMs, Ernie and Qwen, and does not include other models or those from Western contexts.
+However, we carried out a qualitative analysis to investigate the robustness of this approach, which yielded results consistent with this assumption.
+Our work is limited to two Chinese LLMs, Ernie and Qwen, and does not include other models, including those from Western contexts.
+Additionally, since both language models are closed-source, we lack access to crucial internal information, such as model parameters, probability distributions, or training data specifics.
+This restriction makes it difficult to fully understand the factors contributing to the prevalence or intensity of negative content in their outputs.
 Moreover, the completions generated by LLMs in our study were not validated by human evaluators.
@@
 Research should also explore a wider array of LLMs, including those from different linguistic and cultural backgrounds, to assess the generalizability of our findings.
-Finally, developing and implementing effective bias mitigation strategies will be essential to address the ethical challenges posed by LLMs, ensuring their responsible use in various applications.
+Finally, developing and implementing effective bias mitigation strategies will be essential to address the ethical challenges posed by LLMs, promoting their responsible use across various application domains.
+Acknowledgments
+During the preparation of this work, the authors used OpenAI ChatGPT in order to proof-check the grammar of some paragraphs and refine the language.
+After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.